• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name jwexler

    Log entry time 07:58:30 on April 03, 2010

    Entry number 315930

    keyword=Shift Summary - Owl shift 04/03/2010

    Shift summary for owl, 04/03.
    
    00:00 - Bob preparing to do optics studies, now that the beam is back.
    00:15 - Repositioned beam to (0,0) on 4A and 4B.  Asked for 35uA for
    compton test.
    00:20 - VDC, upper scint., and s0 on in both arms for optics study.
    00:30 - Begin optics study. Go to 1uA on tantalum target, then take runs
    for different values of p0, for p0 = 1.063, 1.068, 1.058, 1.053, 1.048,
    1.043.  Runs are 27240-27246, 6558-6564.
    
    03:10 - Access to remove sieves.
    03:25 - Exit access, start long beam recovery plan for Pb target.
    
    04:45 - Have 10nA beam on Pb target #3, being Ahmed & Rupesh's 'lead
    counting' run plan.  Ahmed takes ~1M events for each of the three
    detector positions from the previous night (listed in entries 315682,
    315688, 315689).
    
    06:15 - Finish lead counting run plan. Call for escorted access to put
    detector into integrating mode, and call Alex to install a GEM per his
    request.
    
    7:30 - Come out of access, complete listed goals and position A-T hole
    detectors at the edge of their range of motion in y.  Begin short beam
    recovery plan on Pb target, begin A-T scan run plan.